OpenStudy (anonymous):

5/13t = -9 solve this and please show the process of how it's the answer. Thank you!

OpenStudy (wolf1728):

5/13t = -9 Multiplying both sides by 13t 5 = -117t Dividing both sides by -117 5/-117 = t -0.0427350427 = t